OpenStudy (anonymous):

Two Bodies are dropped from two different heights h1 and h2 one of them is dropped directly down while the other is projected with a horizontal velocity of 10m?s It is found that both reach the ground with same speed (g=10m/s^2) Then h1/h2 = root10 h1-h2=5m

OpenStudy (anonymous):

FOR BODY DROPPED DIRECTLY. 2ah1 = v^2 => v = sqrt(2gh1)----- = speed----2 For BODY 2. Vertical velocity is sqrt ( 2gh2) Horizontal is 10. Therfore speed = sqrt ( 2*10*h2 + 100) = sqrt(20h2 + 100) ----1 Equating 1 and 2, 20h1 = 20h2 + 100 h1 - h2 = 5

OpenStudy (anonymous):

@siddhantsharan i did nt understand 2 nd eq

OpenStudy (anonymous):

Therfore speed = sqrt ( 2*10*h2 + 100) = sqrt(20h2 + 100)???????????

OpenStudy (anonymous):

The vertical component will be \[\sqrt{2*g*h2}\] Horizontal is 10m/s. Now these are two vectors. Therefore Their net resultant is \[\sqrt{a^2 + b^2}\] Where a and b are the vertical and horizontal velocities,

OpenStudy (anonymous):

@siddhantsharan ok then how eq 1 = eq2

OpenStudy (anonymous):

Its given that the speeds are equal.

OpenStudy (anonymous):

ok sorry i did nt see it thanzzzzz a lot

OpenStudy (anonymous):

@siddhantsharan Vertical velocity is sqrt ( 2gh2) how???

OpenStudy (anonymous):

\[v^2 - u^2 = 2as\] You know that. Now, s = h2 and u = 0 Therefore, \[v^2 = 2hg\]